OpenStudy (anonymous):

Let Vx and Vy be the volumes of the solids that result when the region enclosed by y=1/x y=0 x=1/2 and x=1/2 (b > 1/2) is revolved about the x-axis and the y-axis, respectively. Is there a value of b for which Vx = Vy?

OpenStudy (anonymous):

@sourwing i accidentally closed my question. would you be able to help ? :/

OpenStudy (anonymous):

well, you see, the question doesn't ask for what it is. It's yes/no question :DDD

OpenStudy (anonymous):

@sourwing hehe i wish it was as simple as that! the answer manual actually gives an actual number so I think i have to support the yes

OpenStudy (anonymous):

you have the solution manual and still don't know the answer? O.O

OpenStudy (anonymous):

Well Idk how they got the answer it only give me the answer

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

it's revolved around the x-axis, what is you setup?

OpenStudy (anonymous):

well this is where I have my doubts. Im not sure where to start!

OpenStudy (anonymous):

V_x = pi ∫ radius^2 [disk method] V_x = pi ∫(1/x)^2 dx, from 1/2 to b around the y-axis: V_y = 2pi ∫(radius) (height) [shell method] V_y = 2pi ∫ x (1/x) dx from 1/2 to b

OpenStudy (anonymous):

set V_x = V_y and solve for b

OpenStudy (anonymous):

oh thank you! let me try it out

OpenStudy (anonymous):

for Vx I got..\[2\pi \left( \frac{ 4b ^{2}-1 }{ b ^{2} } \right)\]

OpenStudy (anonymous):

and my Vy=2pi

OpenStudy (anonymous):

wait for the Vx just pi not 2pi!

OpenStudy (dan815):

can you exaplain ur question to me

OpenStudy (dan815):

why are there 2, x=1/2s

OpenStudy (anonymous):

oops I'm sorry i may have wrote that twice it is supposed to say…

OpenStudy (dan815):